2022 2021 Net sales $4,750.0 $5,500.6 Cost of goods sold 3,400.4 3,900.8 Net income 65.1 150.5 Accounts receivable 75.1 107.7 Inventory 1,150.0 1,250.0 Total assets 2,950.1 3,250.1 Total common stockholders’ equity 900.2 1,130.8 Compute the following ratios for 2022. i need help with the asset turnover please?
Answer #1

Please find the following solution for Asset Turnover Ratio.

Net sales for 2022 = $ 4,750.00

Total assets 2022 = $ 2950.1

Total assets 2021 = $ 3250.1

Asset Turnover Ratio = Net Sales / Average Total Asset

= 4,750 / ( 2,950.1 + 3,250.1 ) / 2

= 4,750 / ( 2,950.1 + 3,250.1 ) / 2

= 4,750 / 3100.1

Asset Turnover Ratio = 1.5322